# On the stability of the penalty function for the $\mathbb{Z}^{2}$-hard square shift

### v2

@@ Abstract.
We investigate the stability of maximizing measures for a penalty function of a two-dimensional subshift of finite type, building on the work of Gonschorowski et al. [GQS]. In the one-dimensional case, such measures remain stable under Lipschitz perturbations for any subshift of finite type. However, instability arises for a penalty function of the Robinson tiling, which is a two-dimensional subshift of finite type with no periodic point and zero entropy. This raises the question of whether stability persists in two-dimensional subshifts of finite type with positive topological entropy. In this paper, we address this question by studying the $\mathbb{Z}^{2}$-hard square shift, a well-known example of a two-dimensional subshift with positive entropy. Our main theorem establishes that, in contrast to previous results, a penalty function of the hard square shift remains stable under Lipschitz perturbations.
@@ 1. Introduction
In [GQS], the authors highlight the difference between one and two dimensions, demonstrating that instability arises in a penalty function of the Robinson tiling, which is a two-dimensional subshift of finite type with no periodic points and zero entropy. This naturally prompts the question of whether stability persists for a two-dimensional subshift of finite type with positive topological entropy. In this paper, we address this question by investigating the $\mathbb{Z}^{2}$-hard square shift, a well-known example of a two-dimensional subshift of finite type with positive entropy (See [Pav12] for more properties of hard square shifts). Our main theorem establishes that, in contrast to the results on $\mathbb{Z}^{2}$ subshift of finite type as presented by Gonschorowski et al., the penalty function of the hard square shift remains stable under Lipschitz perturbations.
Informally, a subshift of finite type is defined by specifying a finite set of finite “forbidden patterns” $F$ made up of letters from alphabet $\mathcal{A}$, and defining $X_{F}$ to be the set of configurations in $\mathcal{A}^{\mathbb{Z}^{d}}$ in which no pattern from $F$ appears (see §2 for more details). The set $F$ is called a forbidden set. The $\mathbb{Z}^{2}$-hard square shift is a subshift of finite type where no two adjacent 1’s appear, either horizontally or vertically. Although other forbidden sets can be used to define the hard square shift, we will use the following forbidden set $F$ to define the penalty function later:
(1) $F=\left\{\begin{smallmatrix}1&1\\ 1&1\end{smallmatrix}~{}~{}\begin{smallmatrix}1&0\\ 1&0\end{smallmatrix}~{}~{}\begin{smallmatrix}1&0\\ 1&1\end{smallmatrix}~{}~{}\begin{smallmatrix}0&0\\ 1&1\end{smallmatrix}~{}~{}\begin{smallmatrix}0&1\\ 1&1\end{smallmatrix}\right\}.$
We now define the penalty function as follows:
$f(x)=\left\{\begin{array}[]{cc}-1&\mbox{if}~{}\begin{smallmatrix}x_{(0,1)}&x_{(1,1)}\\ x_{(0,0)}&x_{(1,0)}\end{smallmatrix}\in F\\ 0&\mbox{otherwise}\end{array}\right..$
@@
Let $X$ be the hard square shift and $f$ be the penalty function. Then there exists $\varepsilon>0$ such that for every Lipschitz continuous function $g$ with $\|f-g\|_{{\rm Lip}}<\varepsilon$, every maximizing measure of $g$ is supported on $X$.
We remark that the stability result for the hard square shift is relatively straightforward, since forbidden words can be easily eliminated by replacing $1$ with 00. However, extending this result to other nearest neighbor subshifts of finite type, such as those discussed in [Pav12], appears to be more challenging. This difficulty arises because, in general, there is no method to identify the exact locations of forbidden words, making it impossible to estimate the distances required to remove them.
@@ 2. Settings
for $\underline{x},\underline{y}\in\mathcal{A}^{\mathbb{Z}^{2}}$ where $i=\inf\{\|u\|_{\infty}:x_{u}\neq y_{u}\}$. Then, this metric is compatible with the product topology.
@@
For any full shift $\mathcal{A}^{\mathbb{Z}^{2}}$, we define the $\mathbb{Z}^{2}$-action $\{\sigma_{u}\}_{u\in\mathbb{Z}^{2}}$ on $\mathcal{A}^{\mathbb{Z}^{2}}$ as follows: for any $u\in\mathbb{Z}^{2}$ and $\underline{x}\in\mathcal{A}^{\mathbb{Z}^{2}}$, $(\sigma_{u}(\underline{x}))_{v}=x_{u+v}$ for all $u\in\mathbb{Z}^{2}$.
A subset $X\subset\mathcal{A}^{\mathbb{Z}^{2}}$ is a subshift if it is closed and shift-invariant, i.e. for any $\underline{x}$ and $u\in\mathbb{Z}^{2}$, $\sigma_{u}(\underline{x})\in X$.
@@
A configuration $w$ on the alphabet $\mathcal{A}$ is any mapping from a non-empty subset $S$ of $\mathbb{Z}^{2}$ to $\mathcal{A}$, where $S$ is called the shape of $w$. For any configuration $w$ with shape $S$ and any $T\subset S$, denote by $w|_{T}$ the restriction of $w$ to $T$, i.e. the subconfiguration of $w$ appearing $T$.
A subshift $X$ is a shift of finite type if there exists a finite collection $F$ of finite configuration on $\mathcal{A}$ such that $X=X_{F}$ to be the set of $\underline{x}\in\mathcal{A}^{\mathbb{Z}^{2}}$ such that $\underline{x}|_{S}\notin F$ for all finite $S\subset\mathbb{Z}^{2}$.
Definition 2.1 (The $\mathbb{Z}^{2}$-hard square shift).
Let $\mathcal{A}=\{0,1\}$ and set the forbidden set by
$F=\left\{\begin{smallmatrix}1&1\\ 1&1\end{smallmatrix}~{}~{}\begin{smallmatrix}1&0\\ 1&0\end{smallmatrix}~{}~{}\begin{smallmatrix}1&0\\ 1&1\end{smallmatrix}~{}~{}\begin{smallmatrix}0&0\\ 1&1\end{smallmatrix}~{}~{}\begin{smallmatrix}0&1\\ 1&1\end{smallmatrix}\right\}.$
@@
Then the subshift $X=X_{F}$ is called the $\mathbb{Z}^{2}$-hard square shift.
It is easy to see that the topological entropy of the hard square shift is positive.
For any $a,b\in\mathbb{Z}$ with $a<b$, we use $[a,b]$ to denote $\{a,a+1,\ldots,b\}$. For each $n\geq 0$ define the box of size $n$ as
$\Lambda_{n}=[-n,n]\times[-n,n].$
@@
The cardinality of $\Lambda_{n}$ is given by $\lambda_{n}=\#\Lambda_{n}=(2n+1)^{2}$. For a continuous function $f$ and a nonempty subset $T\subset\mathbb{Z}^{2}$ define a dynamical sum over $T$ by
$S_{T}f=\sum_{u\in T}f\circ\sigma^{u}.$
@@ 3. Proof of the main theorem
Lemma 3.1 ([GQS, Lemma 2.1.]).
Let $J\subset X$ be a subset of a compact metric space $X$ and $f$ be a Lipschitz continuous function with $f|_{J}=0$. For $\varepsilon>0$ and a Lipschitz continuous function $g$ with $\|f-g\|_{{\rm Lip}}<\varepsilon$ we have
@@
This lemma will be applied in our setting with $B=f^{-1}\{0\}$. With this preparation, we now proceed to the proof of our main theorem. A key feature of this proof is its extension of the coupling and splicing argument, as well as the ”path-wise surgery” technique from [GQS], to a two-dimensional case.
@@
Let $\underline{x}$ be a generic point for the measure $\mu$. For each $i\in\mathbb{Z}$ let
$S^{i}=\{u_{1}\in\mathbb{Z}:\sigma^{(u_{1},i)}\underline{x}\in I^{c}\}.$
For $u_{1},v_{1}\in S^{i}$ set a relation $u_{1}rv_{1}$ by
$u_{1}rv_{1}\Leftrightarrow|u_{1}-v_{1}|=1.$
Moreover define the equivalent relation $\sim$ on $S^{i}$ by
@@
$u_{1}\sim v_{1}\Leftrightarrow\ \mbox{there exist}\ w_{1}^{1},\ldots,w_{1}^{p}\in S^{i}\ \mbox{such that}\ u_{1}rw_{1}^{1};w_{1}^{1}rw_{1}^{2};\cdots;w_{1}^{p}rv_{1}.$
Then we get the sequence of bad words on ×$\mathbb{Z}\times\{i\}$. Set $S^{i}/\sim=\{B_{n}^{i}\}_{n\in\mathbb{Z}}$ where $B_{0}^{i}$ is the equivalent class including 00 if $0\in S^{i}$and it is the equivalent class including $\min\{u_{1}>0:(u_{1},i)\in S^{i}\}$ if $0\notin S$. Setting $\alpha^{i}_{n}=\min\{u_{1}\in B_{n}^{i}\}$ and $\beta_{n}^{i}=\max\{u_{1}\in B_{n}^{i}\}$ for each $n$, we have $B_{n}^{i}=[\alpha^{i}_{n},\beta^{i}_{n}]$.
@@
Then we define a sequence of configurations by inductively replacing $1$ to 00 on each line. First set, $\underline{x}^{(-1)}:=\underline{x}$ and define $\underline{x}^{(0)}$ by
$x^{(0)}_{(u_{1},u_{2})}=\left\{\begin{array}[]{cc}0&\mbox{if}\ u_{1}\in S^{0}\ \mbox{and}\ u_{2}=0\\ x_{(u_{1},u_{2})}&\mbox{othewise}.\end{array}\right.$
For $k\geq 1$, we define $\underline{x}^{(2k-1)}$ and $\underline{x}^{(2k)}$ inductively as follows:
$x^{(2k-1)}_{(u_{1},u_{2})}$ =$=\left\{\begin{array}[]{cc}0&\mbox{if}\ u_{1}\in S^{k}\ \mbox{and}\ u_{2}=k\\ x^{2k}_{(u_{1},u_{2})}&\mbox{otherwise},\end{array}\right.$
$x^{(2k)}_{(u_{1},u_{2})}$ =$=\left\{\begin{array}[]{cc}0&\mbox{if}\ u_{1}\in S^{-k}\ \mbox{and}\ u_{2}=-k\\ x^{2k-1}_{(u_{1},u_{2})}&\mbox{otherwise}.\end{array}\right.$
Then there is no bad word in $\underline{x}^{(2k-1)}$ on $[-k+1,k]\times\mathbb{Z}$ and in $\underline{x}^{(2k)}$ on $[-k,k]\times\mathbb{Z}$.
Furthermore, define $\tilde{x}$ by
$\tilde{x}_{(u_{1},u_{2})}=\left\{\begin{array}[]{cc}0&\mbox{if}\ u_{1}\in S^{k},u_{2}=k\ \mbox{and}\ k\in\mathbb{Z}\\ x_{(u_{1},u_{2})}&\mbox{otherwise}.\end{array}\right.$
Clearly, $\tilde{x}\in X$. By definition, we have $\lim_{k\to\infty}\underline{x}^{(k)}=\tilde{x}$.
@@
Fix sufficiently large $N\geq 1$. We now consider the difference between the dynamical sums of $\underline{x}$ and $\tilde{\underline{x}}$ over $\Lambda_{N}$ :
(2) $S_{\Lambda_{N}}g(\underline{x})-S_{\Lambda_{N}}g(\tilde{\underline{x}})$ $=\sum_{i=0}^{2N-1}\left(S_{\Lambda_{N}}g(\underline{x}^{i-1})-S_{\Lambda_{N}}g(\underline{x}^{i})\right)+S_{\Lambda_{N}}g(\underline{x}^{(2N)})-S_{\Lambda_{N}}g(\tilde{\underline{x}}).$
@@
(3) $S_{\Lambda_{N}}g(\underline{x}^{(2N)})-S_{\Lambda_{N}}g(\tilde{\underline{x}})$ $=2\sum_{i=1}^{N}\frac{(2N+1)}{2^{i}}\leq 2(2N+1).$
In order to provide an upper bound for the summation term, we analyze the difference between the dynamical sums of $\underline{x}^{i-1}$ and $\underline{x}^{i}$ over $\Lambda_{N}$ by considering contributions from bad words and good words separately.
Estimate on bad words: Let $i$ be a nonnegative integer. For $n\in\mathbb{Z}$ and $u_{1}\in B_{n}^{i}$ we have $g(\sigma^{(u_{1},i)}\underline{x}^{(i-1)})<-1+\varepsilon$ and $g(\sigma^{(u_{1},i)}\underline{x}^{(i)})>-\varepsilon$. Hence we have
$S_{B_{n}^{i}\times\{i\}}g(\underline{x}^{i-1})-S_{B_{n}^{i}\times\{i\}}g(\underline{x}^{i})<|B_{n}^{i}|(-1+\varepsilon)+|B_{n}^{i}|\varepsilon=|B_{n}^{i}|(-1+2\varepsilon)$
@@
Estimate on Good words: To establish an estimate for good words, we introduce the following notations: For a nonnegative integer $i$ set $I^{i}_{N}$ by
$I^{i}_{N}=\{n:B_{n}^{i}\cap[-N,N]\}$
and let $\ell_{N}^{i}=\min I_{n}^{i}$ and $r_{N}^{i}=\max I_{n}^{i}$. Since we are interested only in bad words within $\Lambda_{N}$, we modify the endpoints as follows. For $n\in[\ell_{N}^{i}-1,r_{N}^{i}+1]$ set
$\tilde{\alpha}_{n}^{i}=\left\{\begin{array}[]{ccc}-N&\mbox{if}&-N\in S^{i}\ \mbox{and}\ n=\ell_{N}^{i}\\ N+1&\mbox{if}&n=r_{N}^{i}+1\\ \alpha_{n}^{i}&\mbox{else}&\end{array}\right.$
and
$\tilde{\beta}_{n}^{i}=\left\{\begin{array}[]{ccc}-N&\mbox{if}&n=\ell_{N}^{i}-1\\ N&\mbox{if}&N\in S^{i}\ \mbox{and}\ n=r_{N}^{i}\\ N+1&\mbox{if}&N\notin S^{i}\ \mbox{and}\ n=r_{N}^{i}+1\\ \beta_{n}^{i}&\mbox{else.}\end{array}\right.$
Refer to Figure 1 for a visual representation of this operation.
[…]
Case1111-i : $-N\in S^{i}$
[…]
Case1111-ii : $-N\notin S^{i}$
[…]
Case2222-i : $N\in S^{i}$
[…]
Case2222-ii : $N\notin S^{i}$
Figure 1. The value of $\tilde{\alpha}_{n}^{i},\tilde{\beta}_{n}^{i}$

### v3

# On the stability of the penalty function for a nearest-neighbor $\mathbb{Z}^{2}$ subshift of finite type with the single-site fillability

@@ Abstract.
We investigate the stability of maximizing measures for a penalty function of a two-dimensional subshift of finite type, building on the work of Gonschorowski et al. [GQS]. In the one-dimensional case, such measures remain stable under Lipschitz perturbations for any subshift of finite type. However, instability arises for a penalty function of the Robinson tiling, which is a two-dimensional subshift of finite type with no periodic points and zero entropy. This raises the question of whether stability persists in two-dimensional subshifts of finite type with positive topological entropy. In this paper, we address this question by studying a nearest-neighbor subshift of finite type satisfying the single-site fillability property. Our main theorem establishes that, in contrast to previous results, a penalty function of such a subshift of finite type remains stable under Lipschitz perturbations.
@@ 1. Introduction
In [GQS], the authors highlight the difference between one and two dimensions, demonstrating that instability arises in the penalty function of the Robinson tiling, a two-dimensional subshift of finite type with no periodic points and zero entropy. In contrast, in the one-dimensional setting, stability results are established for subshifts of finite type that typically possess abundant periodic points and positive topological entropy. This contrast raises the natural question of whether stronger topological properties—such as positive topological entropy, rich periodic structure, or some mixing property—might guarantee stability in the two-dimensional case as well. In this paper, we consider this question by investigating the penalty function on nearest-neighbor $\mathbb{Z}^{2}$ subshifts of finite type (n.n. SFTs) that satisfy the single-site fillability (SSF) property (See for definitions in §2). This class includes, as a typical example, the hard square shift, a well-known two-dimensional subshift of finite type with positive entropy. Our main theorem establishes that, in contrast to the result on $\mathbb{Z}^{2}$ subshifts of finite type presented by Gonschorowski et al., the penalty function of any such system remains stable under Lipschitz perturbations.
Informally, a subshift of finite type is defined by specifying a finite set of finite “forbidden patterns” $F$ made up of letters from an alphabet $\mathcal{A}$, and defining $X_{F}$ to be the set of configurations in $\mathcal{A}^{\mathbb{Z}^{2}}$ in which no pattern from $F$ appears (see §2 for more details). The set $F$ is called a forbidden set. A subshift of finite type is called a nearest-neighbor subshift of finite tyep (n.n. SFT) if $F$ can be chosen to consist only of patterns supported on pairs of adjacent sites. For a n.n. SFT with a forbidden set $F$ we define the penalty function as follows:
$f(x)=\left\{\begin{array}[]{cc}-1&\mbox{if}\quad x_{(0,0)}x_{(0,1)}\in F\quad\mbox{or}\quad x_{(0,0)}x_{(1,0)}\in F,\\ 0&\mbox{otherwise}.\end{array}\right..$
@@
Let $X$ be a $\mathbb{Z}^{2}$ nearest-neighbor subshift of finite type with the single-site fillability and $f$ be the penalty function. Then there exists $\varepsilon>0$ such that for every Lipschitz continuous function $g$ with $\|f-g\|_{{\rm Lip}}<\varepsilon$, every maximizing measure of $g$ is supported on $X$.
We remark that the stability result for a n.n. SFT satisfying SSF is relatively straightforward since forbidden patterns can be easily eliminated by making local modifications guaranteed by SSF. However, extending this result to more general SFTs appears to be substantially more difficult, as in such problems there is no method to extract the precise locations of bad words, and only their proportion can be accessed. This lack of positional information prevents us from effectively handling configurations. As a result, extending the stability result to systems with properties such as block gluing would likely require fundamentally new techniques.
@@ 2. Settings
We denote the origin $(0,0)$ of $\mathbb{Z}^{2}$ by $\bm{0}$ to simplify notation. For $\bm{u},\bm{v}\in\mathbb{Z}^{d}$ are said to be adjacent if $|\bm{u}-\bm{v}|=1$, where $|\bm{u}|=|u_{1}|+|u_{2}|$ for $\bm{u}=(u_{1},u_{2})\in\mathbb{Z}^{2}$. The boundary of a set $S\subset\mathbb{Z}^{2}$, denoted by $\partial S$, is the set of $\bm{v}\in\mathbb{Z}^{2}\setminus S$ which are adjacent to some element of $S$. For any $a,b\in\mathbb{Z}$ with $a<b$, we use $[a,b]$ to denote $\{a,a+1,\ldots,b\}$. For each $n\geq 0$ define the box of size $n$ as
$\Lambda_{n}=[-n,n]\times[-n,n].$
The cardinality of $\Lambda_{n}$ is given by $\lambda_{n}=\#\Lambda_{n}=(2n+1)^{2}$.
@@
for $\underline{x},\underline{y}\in\mathcal{A}^{\mathbb{Z}^{2}}$ where $i=\inf\{\|\bm{u}\|_{\infty}:x_{\bm{u}}\neq y_{\bm{u}}\}$. Then, this metric is compatible with the product topology.
@@
For any full shift $\mathcal{A}^{\mathbb{Z}^{2}}$, we define the $\mathbb{Z}^{2}$-action $\{\sigma_{\bm{u}}\}_{\bm{u}\in\mathbb{Z}^{2}}$ on $\mathcal{A}^{\mathbb{Z}^{2}}$ as follows: for any $\bm{u}\in\mathbb{Z}^{2}$ and $\underline{x}\in\mathcal{A}^{\mathbb{Z}^{2}}$, $(\sigma^{\bm{u}}(\underline{x}))_{\bm{v}}=x_{\bm{u}+\bm{v}}$ for all $\bm{u}\in\mathbb{Z}^{2}$.
@@
A configuration $w$ on the alphabet $\mathcal{A}$ is any mapping from a non-empty subset $S$ of $\mathbb{Z}^{2}$ to $\mathcal{A}$, where $S$ is called the shape of $w$. If $S$ is finite, we call a configuration $w$ on $S$ is finite. For any configuration $w$ with shape $S$ and any $T\subset S$, we denote by $w|_{T}$ the restriction of $w$ to $T$, i.e., the subconfiguration of $w$ supported on $T$. Let $S,T\subset\mathbb{Z}^{2}$ with $S\cap T=\emptyset$, and let $w$ and $w^{\prime}$ be configurations with shapes $S$ and $T$, respectively. Then the concatenation of $w$ and $w^{\prime}$ is the configuration on $S\cup T$ defined by $(ww^{\prime})|_{S}=w$ and $(ww^{\prime})|_{T}=w^{\prime}$, which is denoted by $ww^{\prime}$. Let $\mathcal{A}^{*}$ be the set of all configurations defined on finite subsets of $\mathbb{Z}^{2}$.
A subset $X\subset\mathcal{A}^{\mathbb{Z}^{2}}$ is a subshift if it is closed and shift-invariant, i.e. for any $\underline{x}$ and $\bm{u}\in\mathbb{Z}^{2}$, $\sigma^{\bm{u}}(\underline{x})\in X$. It is well known that any subshift can be also defined in terms of forbidden patterns: for a countable family $F$ of finite configurations, define
$X=X_{F}=\{\underline{x}\in\mathcal{A}^{\mathbb{Z}^{2}}\mid\sigma^{\bm{u}}(\underline{x})|_{S}\notin F\ \mbox{for all finite}\ S\subset\mathbb{Z}^{d},\ \mbox{for all}\ \bm{u}\in\mathbb{Z}^{2}\}.$
Then $X=X_{F}$ is a subshift and all subshift can be represented in this way.
A subshift $X$ is a shift of finite type (SFT) if there exists a finite collection $F\subset\mathcal{A}^{*}$, called the forbidden set, such that $X=X_{F}$. If $F$ consists only of configurations on pairs of adjacent sites, $X$ is called a nearest-neighbor shift of finite type (n.n. SFT). Hereafter, for nearest-neighbor SFTs, we assume without further comment that their forbidden sets consist only of patterns defined on shapes of the form $\{\bm{0},\bm{0}+e_{i}\}$ for $i=1,2$.
Let $X$ be a subshift. A configuration $w$ on a $S\subset\mathbb{Z}^{2}$ is globally admissible for $X$ if there exists $\underline{x}\in X$ such that $x|_{S}=w$. Let $F$ be a forbidden set defining $X$. A configuration $w$ on $S\subset\mathbb{Z}^{2}$ is locally admissible for $X=X_{F}$ if for every $S^{\prime}\subset S$, $w|_{S}\neq F$, up to translation.
Many combinatorial and topological mixing properties have been studied in [MP15, Br16, BMP18]. Here, we consider a strong combinatorial mixing property, the single-site fillability, introduced in [MP15].
Definition 2.1.
A n.n. SFT $X$ is single-site fillable (SSF) if for a finite forbidden set $F\subset\mathcal{A}^{*}$ such that $X=X_{F}$ and for every configuration $w$ on the shape on $\mathcal{A}^{\partial\{\bm{u}\}}$ for some $\bm{u}\in\mathbb{Z}^{2}$, there exists $a\in\mathcal{A}^{\{\bm{u}\}}$ such that $wa$ is locally admissible.
The property SSF is a generalization of the concept of a safe symbol. The symbol $a\in\mathcal{A}^{\{\bm{0}\}}$ in the definition of SSF may depend on the configuration $w\in\mathcal{A}^{\partial\{\bm{0}\}}$. If such a symbol $a$ can be chosen independently of the surrounding configuration $w$, then $a$ is called a safe symbol (see for example [MP15, Br16] for discussions on the relationships between mixing properties). Note that for n.n. SFT $X=X_{F}$, a locally admissible configuration is globally admissible[MP15].
The following are typical examples of n.n. SFTs.
Example 2.2 (Hard square shift).
Let $\mathcal{A}=\{0,1\}$. Define $F\subset\mathcal{A}^{*}$ by
$F=\bigcup_{i=1,2}\{w:\{\bm{0},\bm{0}+e_{i}\}\rightarrow\mathcal{A}\mid w(\bm{0})=w(\bm{0}+e_{i})=\{1\}\}.$
@@
Then $X_{F}$ is called the hard square shift, which consists of configurations $\underline{x}\in\mathcal{A}^{\mathbb{Z}^{2}}$ with no adjacent 1’s.
Example 2.3 ($k$-Checkerboard shift).
Let $k\geq 2$, $\mathcal{A}=\{0,1,\ldots,k-1\}$. Define $F$ by
$F=\bigcup_{i=1,2}\{w:\{\bm{0},\bm{0}+e_{i}\}\rightarrow\mathcal{A}\mid w(\bm{0})=w(\bm{0}+e_{i})\}.$
Then $X_{F}$ is called the $k$-checkerboard shift, which consists of configurations $\underline{x}\in\mathcal{A}^{\mathbb{Z}^{2}}$ where no two adjacent sites have the same symbol.
The hard square shift has a safe symbol 00, as replacing any letter in a configuration by 00 yields an admissible configuration. In contrast, the $k$-checkerboard shift does not have a safe symbol for any $k\geq 2$, since changing a letter arbitrarily may create adjacent sites with the same symbol. However, for $k\geq 5$, the $k$-checkerboard shift satisfies SSF, since there are always enough remaining symbols to fill a site without violating adjacency constraints.
@@
For a continuous function $f$ and a nonempty subset $T\subset\mathbb{Z}^{2}$ define a Birkhoff sum over $T$ by
$S_{T}f=\sum_{\bm{u}\in T}f\circ\sigma^{\bm{u}}.$
@@ 3. Proof of the main theorem
Lemma 3.1 (A version of [GQS, Lemma 2.1.]).
Let $J\subset X$ be a subset of a compact metric space $X$ and $f$ be a Lipschitz continuous function with $f|_{J}=c$ for some constant $c\in\mathbb{R}$. For $\varepsilon>0$ and a Lipschitz continuous function $g$ with $\|f-g\|_{{\rm Lip}}<\varepsilon$ we have
@@
This lemma will be applied in our setting with $J=f^{-1}\{0\}$ and also with $J=f^{-1}\{-1\}$. With this preparation, we now proceed to the proof of our main theorem. A key feature of this proof is its extension of the coupling and splicing argument, as well as the ”path-wise surgery” technique from [GQS], to a two-dimensional case.
@@
Let $\underline{x}$ be a generic point for the measure $\mu$. Let $S^{0}=\{\bm{0}\}$ if $\underline{x}\in I^{c}$, $S^{0}=\emptyset$ otherwise. For each $i\in\mathbb{N}$ let
$S^{i}=\{\bm{u}\in\Lambda_{i}\setminus\Lambda_{i-1}\mid\sigma^{\bm{u}}\underline{x}\in I^{c}\}=S^{t(i)}\sqcup S^{b(i)}\sqcup S^{r(i)}\sqcup S^{l(i)}$
where
$S^{t(i)}$ =$=\{(u_{1},u_{2})\in S^{i}\mid u_{2}=i\},\quad S^{b(i)}=\{(u_{1},u_{2})\in S^{i}\mid u_{2}=-i\},$
$S^{r(i)}$ =$=\{(u_{1},u_{2})\in S^{i}\mid u_{1}=i,u_{2}\notin\{i,-i\}\},$
$S^{l(i)}$ =$=\{(u_{1},u_{2})\in S^{i}\mid u_{1}=-i,u_{2}\notin\{i,-i\}\}.$
Figure 1.
For each $\tau\in\{t(i),b(i),r(i),l(i)\}$ and $\bm{u},\bm{v}\in S^{\tau}$ set a relation $\bm{u}r\bm{v}$ by if they are adjacent, that is,
$\bm{u}r\bm{v}\Leftrightarrow|\bm{u}-\bm{v}|=1.$
Moreover define the equivalent relation $\sim$ on $S^{\tau}$ by
@@
$\bm{u}\sim\bm{v}\Leftrightarrow\ \mbox{there exist}\ \bm{w}^{1},\ldots,\bm{w}^{p}\in S^{\tau}\ \mbox{such that}\ \bm{u}r\bm{w}^{1};\bm{w}^{1}r\bm{w}^{2};\cdots;\bm{w}^{p}r\bm{v}.$
Then we get the segments of bad words on $[-i,i]\times\{i\}$. Set $S^{\tau}/\sim=\{B_{k}^{\tau}\}_{k=1}^{K_{\tau}}$ where the indices $k$ increase from left to right on the top and bottom sides, and from bottom to top on the right and left sides. Setting
$\alpha_{k}^{\tau}=\begin{cases}\min\{u_{1}\mid(u_{1},u_{2})\in B_{k}^{\tau}\}&\text{if }\tau\in\{t(i),b(i)\},\\ \min\{u_{2}\mid(u_{1},u_{2})\in B_{k}^{\tau}\}&\text{if }\tau\in\{r(i),l(i)\},\end{cases}$
and
$\beta_{k}^{\tau}=\begin{cases}\max\{u_{1}\mid(u_{1},u_{2})\in B_{k}^{\tau}\}&\text{if }\tau\in\{t(i),b(i)\},\\ \max\{u_{2}\mid(u_{1},u_{2})\in B_{k}^{\tau}\}&\text{if }\tau\in\{r(i),l(i)\},\end{cases}$
we have
$B_{k}^{\tau}=\begin{cases}[\alpha_{k}^{\tau},\beta_{k}^{\tau}]\times\{i\}&\text{if }\tau=t(i),\\ [\alpha_{k}^{\tau},\beta_{k}^{\tau}]\times\{-i\}&\text{if }\tau=b(i),\\ \{i\}\times[\alpha_{k}^{\tau},\beta_{k}^{\tau}]&\text{if }\tau=r(i),\\ \{-i\}\times[\alpha_{k}^{\tau},\beta_{k}^{\tau}]&\text{if }\tau=l(i).\end{cases}$
Before proving the main theorem, we show the following lemma, which will be used to apply the SSF property.
Lemma 3.2.
For $i\in\mathbb{Z}$, an interval $[\alpha,\beta]\subset\mathbb{Z}$, and a configuration $z$ on $\partial([\alpha,\beta]\times\{i\})\cup([\alpha,\beta]\times\{i\})$, there exists a configuration $w$ on $[\alpha,\beta]\times\{i\}$ such that $z|_{\partial([\alpha,\beta]\times\{i\})}w$ is locally admissible.
The same holds for a configuration on $z$ on ($\partial(\{i\}\times[\alpha,\beta])\cup(\{i\}\times[\alpha,\beta])$.
Proof.
Using SSF inductively, we construct $w$ as follows. First, replace $z|_{\{(\alpha,i)\}}$ by a symbol $a$ such that $z|_{\partial{\{(\alpha,i)\}}}a$ is admissible. Let $z^{(0)}$ denote the resulting configuration on $\partial([\alpha,\beta]\times\{i\})\cup([\alpha,\beta]\times\{i\})$.
Then, for each $j\in\{1,2,\ldots,\beta-\alpha\}$, replace $z^{(j-1)}|_{\{(\alpha+j,i)\}}$ by a symbol $a$ such that $z^{(j-1)}|_{\partial{\{(\alpha+j,i)\}}}a$ is admissible. Define $z^{(j)}$ as the configuration obtained after this replacement.
Finally, set
$w=z^{(0)}_{(\alpha,i)}z^{(1)}_{(\alpha+1,i)}\cdots z^{(\beta-\alpha)}_{(\beta,i)}.$
This $w$ satisfies the desired property. ∎
@@
Then we define a sequence of configurations on $\mathbb{Z}^{2}$ inductively as follows. First, set
$\underline{x}^{(-1)}:=\underline{x}.$
Next define $\underline{x}^{(i)}$ inductively by
$x_{\bm{u}}^{(i)}=x_{\bm{u}}^{(i-1)}\quad\mbox{if}\quad\bm{u}\notin S^{i}$
and replace the configuration on $S^{i}$ by using Lemma 3.2.
Then, by definition, there is no bad word in the configuration $\underline{x}^{(k)}|_{\Lambda_{k}}$ and the sequence $\{\underline{x}^{(k)}\}_{k=0}^{\infty}$ converges. Denote by the limit $\tilde{x}$, and it is clear that $\tilde{x}\in X$.
@@
Fix sufficiently large $N\geq 1$. We now consider the difference between the Birkhoff sums of $\underline{x}$ and $\tilde{\underline{x}}$ over $\Lambda_{N}$ :
(1) $S_{\Lambda_{N}}g(\underline{x})-S_{\Lambda_{N}}g(\tilde{\underline{x}})$ $=\sum_{i=0}^{N}\left(S_{\Lambda_{N}}g(\underline{x}^{i-1})-S_{\Lambda_{N}}g(\underline{x}^{i})\right)+S_{\Lambda_{N}}g(\underline{x}^{(N)})-S_{\Lambda_{N}}g(\tilde{\underline{x}}).$
@@
(2) $S_{\Lambda_{N}}g(\underline{x}^{(N)})-S_{\Lambda_{N}}g(\tilde{\underline{x}})$ $=2\sum_{i=1}^{N}\frac{\#(\Lambda_{N+1}\setminus\Lambda_{N})}{2^{i}}\leq 2(8N+1).$
In order to provide an upper bound for the summation term, we analyze the difference between the Birkoff sums of $\underline{x}^{i-1}$ and $\underline{x}^{i}$ over $\Lambda_{N}$ by considering contributions from bad words and others separately.
Fix $i\geq 0$. First we divide $\Lambda_{N}$ into four regions $\Lambda_{N}=R^{t(i)}\sqcup R^{b(i)}\sqcup R^{r(i)}\sqcup R^{l(i)}$ such that
$R^{t(i)}$ $=[-N,N]\times[i,N]\cup\{(u_{1},u_{2})\mid u_{1}\in[-i,i],|u_{1}|\leq u_{2}\leq i-1\}$
$R^{b(i)}$ $=[-N,N]\times[-N,-i]\cup\{(u_{1},u_{2})\mid u_{1}\in[-i,i],-|u_{1}|\geq u_{2}\geq-i+1,(u_{1},u_{2})\neq\bm{0}\}$
$R^{r(i)}$ $=[i,N]\times[-i+1,i-1]\cup\{(u_{1},u_{2})\mid u_{2}\in[-i+1,i-1],|u_{2}|<u_{1}\leq i-1\}$
$R^{l(i)}$ $=[-i,N]\times[-i+1,i-1]\cup\{(u_{1},u_{2})\mid u_{2}\in[-i+1,i-1],-|u_{2}|>u_{1}\geq-i+1\},$
(see also Figure 2).
Figure 2. The square $ABCD$ represents $\Lambda_{i}$, the square $EFGH$ represents $\Lambda_{N}$, and the polygon $EIAJBKH$ represents $R^{t(i)}$.
Estimate on bad words: For each $\tau\in\{t(i),b(i),r(i),l(i)\}$ and $\bm{u}\in B^{\tau}_{k}$ for some $1\leq k\leq K_{\tau}$, we have
$g(\sigma^{\bm{u}}\underline{x}^{(i-1)})<-1+\varepsilon,\quad\mbox{and}\quad g(\sigma^{\bm{u}}\underline{x}^{(i)})>-\varepsilon.$
Hence we have
$S_{B_{k}^{\tau}}g(\underline{x}^{i-1})-S_{B_{k}^{\tau}}g(\underline{x}^{i})<|B_{k}^{\tau}|(-1+\varepsilon)+|B_{k}^{\tau}|\varepsilon=|B_{k}^{\tau}|(-1+2\varepsilon)$
@@
Estimate on unchanged words: